Suppose that the negative charge in a copper one-cent coin were removed to a very large distance from Earth-perhaps to a distant galaxy-and that the positive charge were distributed uniformly over Earth's surface. By how much would the electric potential at the surface change? (See Sample Problem 22-7 in the problem supplement. By taking the mass of a penny to be 3.11 g, the positive or negative charge is found to be 1.37×105 C)
Answer #1

here,

mass of penny, m = 3.11 g = 0.00311 kg
radius of earth, r = 6371000 m
charge on earth surface, q = 1.37 * 10^5 C

The electric field inside Earth is zero, so the potential remains constant at the value it reaches at the surface:

V = k*q/r
V = (9*10^9 * 1.37*10^5)/(6371000)
V = 1.935*10^8 V